# Optimising the Processing and Storage of Visibilities using lossy compression

### v1

@@ Abstract
We studied the global and local impacts on the output images for continuum and spectral line examples. We found relative error bounds of as much as $10\%$, which provide compression ratios of about 20, have a limited impact on the continuum imaging as the increased noise is less than the image RMS, whereas a $1\%$ error bound (compression ratio of 8) introduces an increase in noise of about an order of magnitude less than the image RMS. For extremely sensitive observations and for very precious data, we would recommend a $0.1\%$ error bound with compression ratios of about 4. These have noise impacts two orders of magnitude less than the image RMS levels. At these levels, the limits are due to instabilities in the deconvolution methods. We compared the results to the alternative compression tool DYSCO, in both the impacts on the images and in the relative flexibility. MGARD provides better compression for similar results, and has a host of potentially powerful additional features.
@@ 1.2 High-performance I/O and Data Compression
Nearly all modern radio astronomy analysis is performed via CASA (casa), which provides an ipython environment and a set of core data processing tasks and utilities. The Casacore Table Data System storage manager can use the Adaptable Input Output System version 2 (ADIOS2) (WANG:2016; Godoy:2020) as the input/output (I/O) and storage backend. ADIOS is a software framework with a simple I/O abstraction and a self-describing data model centred around distributed data arrays, allowing multiple applications to publish and subscribe data at large levels of concurrency. It is primarily focused on high-performance, parallel I/O, with its parallel storage performance, the file format, the memory management, and data aggregation algorithms being designed together to be highly scalable in every axis (many processes, many variables, large amounts of data, many output steps). For the ADIOS/MGARD compression of the DINGO uv-gridded data (williamson-24) we found a seven-fold reduction in the storage footprint and a seven-fold improvement on the processing speed. ADIOS2 additionally provides users the access to state-of-the-art lossless and lossy compressors (i.e. data is fully recovered after decompression, or the data is only approximately recovered) through its operator. By attaching the operator to a variable, ADIOS seamlessly implements the compression and I/O as a combined operation. One example of this is the use of the MGARD library of functions, which we are currently testing as an extension of the software for the SKA pathfinder ASKAP, ASKAPSoft (askapsoft).
MGARD (Gong:2023) is a software that offers error-controlled lossy compression rooted in multi-grid theories. It transforms floating-point scientific data into a set of multilevel coefficients through multi-linear interpolation and $L^{2}$ projection, followed by linear quantisation and lossless encoding processes. The magnitude of the transformed coefficients is close to zero, making them more amenable to compression than the original data. One of MGARD’s notable features is its array of error control options, including the various euclidean norms of $L^{\infty}$, $L^{2}$, point-wise relative $L^{\infty}$, and the ability to vary error bounds across regions or different frequency components. This flexibility is valuable for preserving Region-of-Interest (RoI) and/or Quantities-of-Interest (QoI) (Gong:2022) derived from the reconstructed data. Though we do not explore the use of RoI and QoI in this paper, compression of astronomy data could be significantly improved in size and quality using these features; this is left for future work. For example, RoIs could be used to implement an optimal Baseline Dependent Averaging (BDA) approach on the visibilities.
@@
The natural point of comparison for MGARD is the bit-reduction compression method DYSCO, a lossy compressor specifically designed for radio astronomical data (dysco). DYSCO normalises the data across different antennas, polarisation, timesteps, and frequencies, ensuring a constant noise variance across the full dataset. It then performs non-linear quantisation followed by customised encoding. Unlike MGARD, DYSCO cannot directly prescribe error bounds; the compression-induced errors can only be confirmed post-factum, and its choice of quantisation bins is subjective to the type of normalisation. Previous literature indicates that the main benefit of using DYSCO is that its compression noise does not exhibit spatial structure. In this paper, we demonstrate that MGARD-compressed data exhibit the same properties as well as the previously mentioned advantages.
@@ 1.3 Radio Astronomy Data
Radio data presents a unique challenge. Much of the data is originating from thermal noise as the portion of the sky containing emission is small, see for example Fig. 1 that shows the simulated sky used in these investigations, which is based on the real GLEAM all sky catalogue. In this figure, only a small fraction of pixels in the image contain emission from astronomical sources, which appear as spatially concentrated regions of high radio emission. Not all astronomical sources are spatially concentrated and with ever improving resolution, what was once a single source can be resolved into spatially extended, diffuse emission. Moreover, some signals, such as the sought-after signal of reionisation from the first stars, will be distributed across the entire image and is hidden in the noise (Liu:2014; Nasirudin:2020).
@@ 2.1 Observational Data
We used the SKA simulator for Radio Interferometry data, OSKAR (oskar), to simulate a clean 1 hour-long dataset based on the SKA-Mid AA2 (64 antenna) configuration (hereafter AA2-Mid) and a single polarisation, with the GLEAM catalogue (gleam-cat) to provide 228 unpolarised sources within a $\sim$3 degree field of view. This represents a realistic complex sky, such as would be expected in real observations. This was done at 1.0GHz over a 300MHz bandwidth, with both 100 and 1000 channels. The GLEAM sky-model had a strongest source with a flux of just under 3 Jy, and a standard deviation of about 1 Jy. The baseline lengths with AA2-Mid range between 20 m and 84 km. In addition we added a further column of data, consisting of the GLEAM sky-model plus pure normal-distributed Gaussian noise for each visibility. To represent the continuum case we added the expected per-baseline noise of 0.14 Jy over the whole bandwidth, which was converted to the noise per visibility by scaling with the square root of the number of channels. Thus the thermal noise is several times greater than the sky signal. To investigate a simpler sky, where the sky signals are not varying so rapidly, we replaced the model with a single compact but slightly resolving source, using a few components at the centre of the field. This simple model had an integrated flux of 3 Jy. This was in order to achieve very high dynamic ranges with the limited uv-coverage of the AA2 configuration. To represent the spectral line case we added the simple model to a single channel of the GLEAM sky-model, scaled up by a factor of ten, so that it dominated the thermal noise. This represented a strong, narrow, maser-like emission dataset. We imaged that channel and a few either side to test whether the compression introduces ‘bleed through’ of apparent emission into other spectral channels.
@@
Our final test was to investigate compression on real data; we selected a single typical example of a LOFAR observation. The data selected was observation ID L686982, heavily averaged (for data volume considerations), targeted on the European Large Area ISO Survey deep field N1 at 16:11:00 +54:57:00, averaged down to 230 channels of width 195kHz around 150MHz and 1 minute integrations over the eight hour long observation. The full scientific observations are published (elais-n1) and although our heavily averaged version would not be suitable for a best-quality scientific image, it was suitable to test the behaviour of real data under compression (albeit with improved per visibility signal to noise). With this real data we found that the data distribution was far from Gaussian, because of Radio Frequency Interference (RFI) and Not-a-Number (NaN) values. We truncated the data range to $\pm$100 Jy with all other values flagged and set to zero. The standard deviation of the remaining data was 4 Jy, albeit with a distribution that had an excess of values closer to zero and a long tail of large values. Nevertheless, we compressed it using MGARD in the same fashion as the simulated data, with our best estimate of suitable error bounds.
@@ 2.2 MGARD Compression of complex visibility data
We have compressed the visibilities using MGARD with relative error bounds (EB) between $6\times 10^{-6}-5\times 10^{-1}$, in steps of approximately a factor of 3. Below this range the compression ratios approach unitary and above it one introduces appreciable distortions into the image. The EB is defined as the global root-mean-square-error normalised by the range of data values. Where the range is positive to negative (as in the visibilities) an EB of 0.5 would represent the whole data span and the data could effectively be compressed to a single value for the whole dataset. In practise, for EBs below 0.1, after MGARD compression the absolute error in the reconstructed data values forms a sharply truncated distribution between zero and a few times the error bound, with 99% of the post-compression values below the requested EB. Note that above 0.1 the simple relationship between the EB and the 99% percentile starts to diverge. We tested the impact of data reordering – in time-ordered and baseline-ordered changing fashion – on compression ratios. Finally, for the real data we used an absolute EB, i.e. bounding the compression incurred error by an absolute value rather than percentage relative to data value range, as the appropriate data range was hard to determine.
@@ 2.3 Comparison of MGARD Compression with DYSCO Compression
The results were compared to similar analysis from data compressed with DYSCO, where the compression is limited to bit reduction. That is 32 bit float numbers were reduced to a representation with a lower quantisation; 2-, 3-, 4-, 6-, 8-, 10-, 12-, 16-bit are possible options. After DYSCO compression the absolute error in the reconstructed data values forms a distribution between zero and a maximum value. The distribution does not truncate as sharply as for MGARD, but does allow us to measure an ‘error bound’ in the cumulative probability function below which 99% of the data is reconstructed sufficiently accurately.
In practise, we found that DYSCO did not support the baseline-ordered data format, so for the relevant tests and comparisons all data was in the default time-ordered format and noise dominated.
@@ 2.4 Image Quality Analysis
All the data was imaged with WSClean (wsclean), with an image size of 8000x8000, a cell size of $\!\!^{\prime\prime}$4, a taper of $\!\!^{\prime\prime}$0 and 10,000 clean iterations. This 1 image size does not capture the full field of view for the GLEAM sky-model, so some components do not appear in the image. However they will make a contribution to the visibilities. The choice of imager is arbitrary, as we compared the results from the compressed data to those from the non-compressed data, rather than the input catalogue. The cleaning parameters are also somewhat arbitrary given our comparison methodology.
@@ 3.1 MGARD Compression of different sky models
Figure 4: The compression ratios achieved when compressing the simulated visibility data with various noise profiles. Shown are the compression results from MGARD for data with a complicated sky and a simple sky model. The x-axis describes the relative error bound provided to MGARD for the compression. The (noise-free) simple sky, as the visibilities change in a smoother fashion, exhibit a greater compression than the (noise-free) complicated sky. Note that the compression ratios for all of the data that are dominated by thermal noise are practically identical and overlap on this plot. The highest compression is achieved when the sky model is simple and noise free and ordered by baseline (black line). The noise-free baseline-ordered complicated sky (purple line) is an order of magnitude lower, representing the impact of the rapidly changing sky signals. Finally the time-ordered simple sky has even lower compression ratios, as the time-ordering hides the simple sky signal from the compression algorithm.
@@
We evaluated the compression ratio obtained on the data in the real-imaginary format for a range of simulated sky models. These were: the complex and realistic sky represented the GLEAM sky-model and a simplified compact model at the phase centre (where the model would change smoothly), both with and without added Gaussian random noise, as shown in Figure 4. This data was reorganised to be baseline ordered (i.e. time-fast), allowing the compression algorithm to ‘see’ the baseline data where sky signals would be smoother. The compression shown used MGARD with relative EBs between $10^{-4}$$5\times 10^{-1}$. Where the Gaussian noise dominates, for complicated, simple or noise-only data (blue, red and green), the compression ratio varies from a factor of 3 at EB of 10-4 to several thousands, with about a factor of 20 at an EB of 10-1. The complicated sky model without noise (purple) compressed several times better than the noise. The simple sky model without noise (black) compressed more than an order of magnitude better, with the compression ratio of about a factor of 800 at an EB of 10-1. If the noise-free data was not reordered, so that it was in the default time-ordered format (gold), the compression ratio fell by more than an order of magnitude, to be only marginally better than noise dominated signals. Thus we conclude in a noise-free situation reordering before compression could have an impact, however, as the there are only a few places in the sky where a single simple source dominates over the noise, this will not be a widely applicable domain. As this would be an expensive and IO intensive operation we would not recommend it in general.
@@ 3.4 Image Quality Analysis
The RMS of the image difference in a pixel to pixel comparison between images formed from compressed and non-compressed data is shown in Fig. 7. MGARD-compressed simulated data dominated by signal (noise-free) is marked with a star and the data dominated by noise (noise-added) is marked with a diamond. The real LOFAR data is marked in solid pink lines with circles. The DYCSO-compressed data is marked in purple with squares. Solid lines indicate images that have been deconvolved with CLEAN and dot-dashed lines indicate where no deconvolution was performed. The RMS of the images made with the non-compressed data is shown as the dotted horizontal lines at the top of the plot in the same colours. The RMS of the difference between non-compressed and compressed data images falls significantly below the image residual RMS, indicating that the added noise from the compression is much less than that in the images of the non-compressed data itself. We note the dominant impact is actually in the deconvolved images, where the decovolution reaches an accuracy limit and improved precision in the compression does not improve the reconstruction. For the complicated sky, where the limited number of antennas limit the possible accuracy to about a dynamic range (DR) of 2,000 we see no improvement below an EB of $10^{-3}$. For the simpler sky model, where we can achieve a DR of 14,000, the reconstruction continues to improve beyond this limit, albeit at a lower rate. The real LOFAR data has results very similar to those from the simulations, particularly if we bear in mind the uncertainties in the conversion of absolute to relative EB for the LOFAR data.
@@
Figure 8: Maximum value of the difference between the image from compressed data and the non-compressed data, with the same data labelling as Fig. 7. The maximum error between the images of compressed and non-compressed data picks up (predominately) the change in side-lobes around the strongest sources. Thus, if the image in CLEANed, the values tend to those of the image RMS. For MGARD, with high compression ($>10\%$), the images of compressed data started to significantly diverge from the non-compressed images, whilst DYSCO continues to perform well.
@@
Figure 9: Two point correlation of the images showing the maximum power at all angular scales ($\ell$ from zero to 200, where zero would be a constant value across the $\sim$1 image and 200 would be the power on 16′′ scales) from the cleaned images made with compressed data, normalised by the power in the image made with non-compressed data. On the left is shown are the results for the complicated GLEAM-model, where the reconstruction accuracy reaches a limit at an error bound of $\sim$1E-3. On the right is shown the result for the same data, but imaged without deconvolution, which does not introduce a limit in the reconstruction accuracy, which continues to improve with EB precision. No peak from introduced image artefacts are detectable, and all contributions are well below the noise levels.

### v2

@@ Abstract
We studied the global and local impacts on the output images for continuum and spectral line examples. We found relative error bounds of as much as $10\%$, which provide compression ratios of about 20, have a limited impact on the continuum imaging as the increased noise is less than the image RMS, whereas a $1\%$ error bound (compression ratio of 8) introduces an increase in noise of about an order of magnitude less than the image RMS. For extremely sensitive observations and for very precious data, we would recommend a $0.1\%$ error bound with compression ratios of about 4. These have noise impacts two orders of magnitude less than the image RMS levels. At these levels, the limits are due to instabilities in the deconvolution methods. We compared the results to the alternative compression tool DYSCO, in both the impacts on the images and in the relative flexibility. MGARD provides better compression for similar error bounds, and has a host of potentially powerful additional features.
@@ 1.2 High-performance I/O and Data Compression
Nearly all modern radio astronomy analysis is performed via the Common Astronomy Software Applications (CASA) software package (casa), which provides an ipython environment and a set of core data processing tasks and utilities. The CASACore Table Data System storage manager can use the Adaptable Input Output System version 2 (ADIOS2) (WANG:2016; Godoy:2020) as the input/output (I/O) and storage backend. ADIOS is a software framework with a simple I/O abstraction and a self-describing data model centred around distributed data arrays, allowing multiple applications to publish and subscribe data at large levels of concurrency. It is primarily focused on high-performance, parallel I/O, with its parallel storage performance, the file format, the memory management, and data aggregation algorithms being designed together to be highly scalable in every axis (many processes, many variables, large amounts of data, many output steps). For the ADIOS/MGARD compression of the Deep Investigation of Neutral Gas Origins (DINGO) uv-gridded data (williamson-24) we found a seven-fold reduction in the storage footprint and a seven-fold improvement on the processing speed. ADIOS2 additionally provides users the access to state-of-the-art lossless and lossy compressors (i.e. data is fully recovered after decompression, or the data is only approximately recovered) through its operator. By attaching the operator to a variable, ADIOS seamlessly implements the compression and I/O as a combined operation. One example of this is the use of the MGARD library of functions, which we are currently testing as an extension of the software for the Australian SKA Pathfinder (ASKAP), ASKAPSoft (askapsoft).
MGARD (Gong:2023) is a software that offers error-controlled lossy compression rooted in multi-grid theories. It transforms floating-point scientific data into a set of multilevel coefficients through multi-linear interpolation and $L^{2}$ projection, followed by linear quantisation and lossless encoding processes. The magnitude of the transformed coefficients is close to zero, making them more amenable to compression than the original data. MGARD has been designed to provide error-controlled rather than fixed-bit compression. It guarantees the compression incurred errors to stay below user-prescribed error bounds, such that the lossy compressed data can be trusted for scientific usage. The resulted compression ratios are data dependent with smoother data being more compressible than noisy data. One of MGARD’s notable features is its array of error control options, including the various euclidean norms of $L^{\infty}$, $L^{2}$, point-wise relative $L^{\infty}$, and the ability to vary error bounds across regions or different frequency components. This flexibility is valuable for preserving Region-of-Interest (RoI) and/or Quantities-of-Interest (QoI) (Gong:2022) derived from the reconstructed data. Though we do not explore the use of RoI and QoI in this paper, compression of astronomy data could be significantly improved in size and quality using these features; this is left for future work. For example, RoIs could be used to implement an optimal Baseline Dependent Averaging (BDA) approach on the visibilities.
@@
The natural point of comparison for MGARD is the bit-reduction compression method DYSCO (DYnamical Statistical COmpression), a lossy compressor specifically designed for radio astronomical data (dysco). DYSCO normalises the data across different antennas, polarisation, timesteps, and frequencies, ensuring a constant noise variance across the full dataset. It then performs non-linear quantisation followed by customised encoding. Unlike MGARD, DYSCO cannot directly prescribe error bounds; the compression-induced errors can only be confirmed post-factum, and its choice of quantisation bins is subjective to the type of normalisation. Previous literature indicates that the main benefit of using DYSCO is that its compression noise does not exhibit spatial structure. In this paper, we demonstrate that MGARD-compressed data exhibit the same properties as well as the previously mentioned advantages.
@@ 1.3 Radio Astronomy Data
Radio data presents a unique challenge. Much of the data is originating from thermal noise as the portion of the sky containing emission is small, see for example Fig. 1 that shows the simulated sky used in these investigations, which is based on the real GaLactic and Extragalactic All-sky MWA (GLEAM) radio survey catalogue. In this figure, only a small fraction of pixels in the image contain emission from astronomical sources, which appear as spatially concentrated regions of high radio emission. Not all astronomical sources are spatially concentrated and with ever improving resolution, what was once a single source can be resolved into spatially extended, diffuse emission. Moreover, some signals, such as the sought-after signal of reionisation from the first stars, will be distributed across the entire image and is hidden in the noise (Liu:2014; Nasirudin:2020).
@@ 2.1 Observational Data
We used the SKA simulator for Radio Interferometry data, OSKAR (oskar), to simulate a clean 1 hour-long dataset based on the SKA-Mid AA2 (64 antenna) configuration [^2] https://gitlab.com/ska-telescope/sdp/ska-sdp-par-model/-/tree/add-AA-layouts/data/layouts(hereafter AA2-Mid) and a single polarisation, with the GLEAM catalogue (gleam-cat) to provide 228 unpolarised sources within a $\sim$3 degree field of view. This represents a realistic complex sky, such as would be expected in real observations. This was done at 1.0GHz over a 300MHz bandwidth, with both 100 and 1000 channels. The GLEAM sky-model had a strongest source with a flux of just under 3 Jy, and a standard deviation of about 1 Jy. The baseline lengths with AA2-Mid range between 20 m and 84 km. In addition we added a further column of data, consisting of the GLEAM sky-model plus pure normal-distributed Gaussian noise for each visibility. To represent the continuum case we added the expected per-baseline noise of 0.14 Jy over the whole bandwidth, which was converted to the noise per visibility by scaling with the square root of the number of channels. Thus the thermal noise is several times greater than the sky signal. To investigate a simpler sky, where the sky signals are not varying so rapidly, we replaced the model with a single compact but slightly resolving source, using a few components at the centre of the field. This simple model had an integrated flux of 3 Jy. This was in order to achieve very high dynamic ranges (greater than a thousand) with the limited uv-coverage of the 64 antenna AA2 configuration. To represent the spectral line case we added the simple model to a single channel of the GLEAM sky-model, scaled up by a factor of ten, so that it dominated the thermal noise. This represented a strong, narrow, maser-like emission dataset. We imaged that channel and a few either side to test whether the compression introduces ‘bleed through’ of apparent emission into other spectral channels.
@@
Our final test was to investigate compression on real data; we selected a single typical example of an observation from the LOw-Frequency ARray (LOFAR). The data selected was observation ID L686982, heavily averaged (for data volume considerations), targeted on the European Large Area ISO Survey deep field N1 at 16:11:00 +54:57:00, averaged down to 230 channels of width 195kHz around 150MHz and 1 minute integrations over the eight hour long observation. The full scientific observations are published (elais-n1) and although our heavily averaged version would not be suitable for a best-quality scientific image, it was suitable to test the behaviour of real data under compression (albeit with improved per visibility signal to noise). With this real data we found that the data distribution was far from Gaussian, because of Radio Frequency Interference (RFI) and Not-a-Number (NaN) values. We truncated the data range to $\pm$100 Jy with all other values flagged and set to zero. The standard deviation of the remaining data was 4 Jy, albeit with a distribution that had an excess of values closer to zero and a long tail of large values. Nevertheless, we compressed it using MGARD in the same fashion as the simulated data, with our best estimate of suitable error bounds.
@@ 2.2 MGARD Compression of complex visibility data
We have compressed the visibilities using MGARD with relative error bounds (EB) between $6\times 10^{-6}-5\times 10^{-1}$, in steps of approximately a factor of 3. Below this range the compression ratios approach unitary and above it one introduces appreciable distortions into the image. The EB is defined as the global root-mean-square-error normalised by the range of data values. Where the range is positive to negative (as in the visibilities) an EB of 0.5 would represent the whole data span and the data could effectively be compressed to a single value for the whole dataset. In practise, for EBs below 0.1, after MGARD compression the absolute error in the reconstructed data values forms a sharply truncated distribution between zero and a few times the error bound, with 99% of the post-compression values below the requested EB. Note that above 0.1 the simple relationship between the EB and the 99% percentile starts to diverge. We tested the impact of data reordering – in time-ordered and baseline-ordered changing fashion – on compression ratios. Finally, for the real data we used an absolute EB, i.e. bounding the compression incurred error by an absolute value rather than percentage relative to data value range.
@@ 2.3 Comparison of MGARD Compression with DYSCO Compression
The results were compared to similar analysis from data compressed with DYSCO, where the compression is limited to bit reduction. That is 32 bit float numbers were reduced to a representation with a lower quantisation; 2-, 3-, 4-, 6-, 8-, 10-, 12-, 16-bit are possible options. After DYSCO compression the absolute error in the reconstructed data values forms a distribution between zero and a maximum value. The distribution does not truncate as sharply as for MGARD, but does allow us to measure an ‘error bound’ in the cumulative probability function below which 99% of the data is reconstructed sufficiently accurately. DYSCO also uses a non-linear quantisation scheme, which means that the rarer numbers are less accurately stored. This is harder to compare directly, but we can test the impacts on the data quality estimates derived from the imaging.
In practise, we found that DYSCO did not support the baseline-ordered data format, so for the relevant tests and comparisons all data was in the default time-ordered format and noise dominated. We note that this places our comparison in the domain where DYSCO was reported to perform best; that is the low SNR domain.
@@ 2.4 Image Quality Analysis
All the data was imaged with WSClean (wsclean), with an image size of 8000$\times$8000 px2, a cell size of $\!\!^{\prime\prime}$4, a taper of $\!\!^{\prime\prime}$0 and 10,000 clean iterations. This 1 image size does not capture the full field of view for the GLEAM sky-model, so some components do not appear in the image. However they will make a contribution to the visibilities. The choice of imager is arbitrary, as we compared the results from the compressed data to those from the non-compressed data, rather than the input catalogue. The cleaning parameters are also somewhat arbitrary given our comparison methodology.
@@ 3.1 MGARD Compression of different sky models
Figure 4: The compression ratios achieved when compressing the simulated visibility data with various noise profiles. Shown are the compression results from MGARD for data with a complicated sky (purple squares) and a simple sky (black circles) model. The x-axis describes the relative error bound provided to MGARD for the compression. For the data that are dominated by thermal noise (line-style symbols, multiple colours) the compression ratios as a function of EB are practically identical and overlap on this plot. The highest compression is achieved when the sky model is simple and noise free and ordered by baseline (black line). The noise-free baseline-ordered complicated sky (purple line) is an order of magnitude lower, representing the impact of the rapidly changing sky signals. Finally the time-ordered data is represented with dotted lines. The noise-free time-ordered data has significantly lower compression ratios than the baseline-ordered data, as the time-ordering hides the sky signal from the compression algorithm. In our simulations the complicated and simple sky in time-order provide very similar compression ratios for the same EB bound, particularly for the less aggressive compression, underlining the importance of presenting the sky signal coherently to the compression algorithm.
@@
We evaluated the compression ratio obtained on the data in the real-imaginary format for a range of simulated sky models. These were: the complex and realistic sky represented the GLEAM sky-model and a simplified compact model at the phase centre (where the model would change smoothly), both with and without added Gaussian random noise, as shown in Figure 4. In a subset of the tests this data was reorganised to be baseline ordered (i.e. time-fast), allowing the compression algorithm to ‘see’ the baseline data where sky signals would be smoother. Otherwise it was retained in time-ordered format; for noise-dominated data, for which every sample is independent, the ordering did not affect the compression ratio. The compression shown used MGARD with relative EBs between $10^{-4}$$5\times 10^{-1}$. Where the Gaussian noise dominates, for complicated, simple or noise-only data (blue, red and green), the compression ratio varies from a factor of 3 at EB of 10-4 to several thousands, with about a factor of 20 at an EB of 10-1. The complicated sky model without noise (purple) compressed several times better than the noise. The simple sky model without noise (black) compressed more than an order of magnitude better, with the compression ratio of about a factor of 800 at an EB of 10-1. If the noise-free data was not reordered, so that it was in the default time-ordered format (dotted lines), the compression ratio fell by more than an order of magnitude, to be only few times better than noise dominated signals. For the noise-dominated signals the ordering or the background signal did not matter and all compression ratios were very similar for a given EB. This is inline with expectations, as smooth data will compress best and baseline ordered data, without noise and with a simple sky model, will present the smoothest data to the compression algorithms. Thus we conclude in a noise-free situation reordering before compression could have an impact, however, as the there are only a few places in the sky where a single simple source dominates over the noise, this will not be a widely applicable domain. As this would be an expensive and I/O intensive operation we would not recommend it in general, however it would be an option for strong and simple sources, such as the calibrator scans.
@@ 3.4 Image Quality Analysis
The RMS of the image difference in a pixel to pixel comparison between images formed from compressed and non-compressed data is shown in Fig. 7. MGARD-compressed simulated data dominated by signal (noise-free) is marked with a star and the data dominated by noise (noise-added) is marked with a diamond. The real LOFAR data is marked in solid pink lines with circles. The DYSCO-compressed data is marked in purple with squares. Solid lines indicate images that have been deconvolved with CLEAN and dot-dashed lines indicate where no deconvolution was performed. The RMS of the images made with the non-compressed data is shown as the dotted horizontal lines at the top of the plot in the same colours. The RMS of the difference between non-compressed and compressed data images falls significantly below the image residual RMS, indicating that the added noise from the compression is much less than that in the images of the non-compressed data itself. We note the dominant impact is actually in the deconvolved images, where the decovolution reaches an accuracy limit and improved precision in the compression does not improve the reconstruction. For the complicated sky, where the limited number of antennas limit the possible accuracy to about a dynamic range (DR) of 2,000 we see no improvement below an EB of $10^{-3}$. For the simpler sky model, where we can achieve a DR of 14,000, the reconstruction continues to improve beyond this limit, albeit at a lower rate. The DYSCO points on this plot are slight to the right of the of the equivalent MGARD points, implying that they achieve the same recovered image quality for a larger relative error; this is presumably, as all other aspects are equal, due to the non-linear quantisation. However these gains (approximately improving the performance by a factor of 1.04) are nullified by the better compression for MGARD. The real LOFAR data has results very similar to those from the simulations, particularly if we bear in mind the uncertainties in the conversion of absolute to relative EB for the LOFAR data.
@@
Figure 8: Maximum value of the difference between the image from compressed data and the non-compressed data, with the same data labelling as Fig. 7, including the dotted lines without symbols indicating the RMS of the residual images, for comparison. The maximum error between the images of compressed and non-compressed data picks up (predominately) the change in side-lobes around the strongest sources. Thus, if the image in CLEANed, the values tend to those of the image RMS. For MGARD, with high compression ($>10\%$), the images of compressed data started to significantly diverge from the non-compressed images, whilst DYSCO continues to perform well. This may, however, be more related to the poorer compression ratio of DYSCO.
@@
Figure 9: Two point correlation of the images showing the maximum power at all angular scales ($\ell$ from zero to 200, where zero would be a constant value across the $\sim$1 image and 200 would be the power on 16′′ scales) from the cleaned images made with compressed data, normalised by the power in the image made with non-compressed data. On the left is shown the results for the complicated GLEAM-model, where the reconstruction accuracy reaches a limit at an error bound of $\sim$1E-3. On the right is shown the result for the same data, but imaged without deconvolution, which does not introduce a limit in the reconstruction accuracy, which continues to improve with EB precision. No peak from introduced image artefacts are detectable, and all contributions are well below the noise levels.